What is the inverse function of
f(x)= 3^(x-1) -2

I thought I would get
x= 3^y-1 -2
then x+2 =3^y-1
then logs but I am confused

Answer by lwsshak3(6491)

What is the inverse function of
f(x)= 3^(x-1) -2
y= 3^(x-1) -2
to find the inverse, interchange x and y, then solve for y
..
x=3^(y-1)-2
x+2=3^(y-1)
take log of both sides
log(x+2)=(y-1)log3=ylog3-log3
log(x+2)+log3=ylog3
Inverse,y^-1=(log(x+2)+log3)/log3